# A Framework for Situating Innovations, Opportunities, and Challenges in Advancing Vertical Systems with Large AI Models

### v1

@@ 1 Introduction
Large artificial intelligence (AI) models have long served as powerful tools for advancing domain-specific research and development. Early examples include the adaptation of language embeddings (like word2vec [1]) for generating disease taxonomies [2], as well as the use of YOLO-based image models  [3] for conducting animal population censuses  [4]. More recently, large models such as GPT-4o  [5] and SAM  [6] have demonstrated more advanced capabilities, catalyzing further innovations in domains as varied as AI tutoring  [7] and digital pathology  [8], all the way to software development  [9]. The success of these models has been propelled by the increasing scale of their architectures—rising from a few million parameters in 2013, to hundreds of millions in 2018, and now surpassing a trillion parameters in the most recent deployments [10]—as well as by the sheer volume of data used for their training [11]. Many widely used benchmarks, both general-purpose and domain-specific, have reached saturation or are rapidly approaching it  [12, 13]. At the same time, frameworks built around these large models (such as Application Programming Interfaces; APIs) have lowered barriers to entry [14], spurring a surge in applications and attracting researchers and practitioners from a broad range of disciplines.
@@
Despite these successes, a decade of adapting large models in diverse verticals has highlighted persistent challenges. For example, research teams have discovered that BERT and CLIP-based models can be brittle to small input variations  [15, 16], large language models (LLMs) often exhibit sensitivity to prompt formatting  [17], and performance of large models can diminish in highly specialized settings [8, 18]. Moreover, AI systems sometimes struggle to effectively support diverse user groups, such as users with a lower Need For Cognition  [19]. These challenges, taken together, pose bottlenecks in the vertical-adoption of the large models. The process of adapting large models to verticals demands coordinated efforts from stakeholders across varied disciplines and continues to be an active area of research. Existing frameworks provide good starting points to address some of these problems at a more granular level—either focusing on one component of developing AI systems or a specific vertical. For instance, Ehsan et al. (2023) [20] present a framework to bridge the gap between social and technical aspects of developing explainable AI systems. On the other hand, focusing on a specific vertical, Trotsyuk et al. (2024)  [21] present a framework to address potential misuse of AI in biomedical research. There is still an unmet need for a more comprehensive framework comprising various components required for vertical adoption of large models and has broad applicability towards generalizing to many verticals.
@@
Our proposed framework consists of 4 layers (see Figure 1), (i) starting with large AI models at the bottom, (ii) vertical-agnostic properties, (iii) vertical adaptation, and (iv) finally, vertical-user intermediaries. The 4 layers represent step-wise modular abstractions involved in developing systems with large models that deliver practical value to their intended users. Drawing on case studies from multiple verticals, we discuss how to situate innovations, challenges, and opportunities within each layer of this framework. We also highlight the inherent dynamism among these layers (i.e., how they interact with and influence one another over time). Finally, drawing on observed trends, we present actionable recommendations that benefit researchers and developers across various verticals and disciplines. For instance, we discuss whether aspects like robustness and privacy are better addressed as vertical-agnostic properties or as a vertical-specific concerns, how scoping feedback from many verticals and intermediary problems can lead to more effective development of newer large models, when domain-specific experts can borrow modeling and interfacing techniques from others while innovating on the domain-specific data curation and evaluation methods, and how interfacing AI systems with users remains ripe with opportunities. Collectively, we believe that adopting our framework will (a) guide optimally placed innovations, (b) highlight potential opportunities, and (c) enable transparent cross-disciplinary dialogue.
Who is the framework for? The framework is designed for interdisciplinary teams who want to adopt large AI models in their respective verticals. For teams focused on a single vertical, the framework decomposes the adoption pipeline into modular components. Across multiple verticals, it provides a holistic view of the broader ecosystem that (a) fosters cross-vertical exchange of innovations, opportunities, and challenges (i.e., what can teams in healthcare learn from teams in education) as well as (b) funnels feedback from many verticals to improve the next iteration of large AI models.
2 A framework for advancing vertical systems with AI models
@@
Large AI models: These are large-scale (in terms of training dataset size and number of parameters) AI models, more recently dubbed foundation models [22] that power diverse applications across verticals. These include modality-specific models (e.g., language-only, vision-only) and multimodal models capable of handling multiple input modalities. Their functional utility lies in the general off-the-shelf capabilities they provide (e.g., natural language understanding, image-text alignment) and adaptability to new verticals through techniques such as fine-tuning, prompting, or in-context learning. Vertical-agnostic properties: Large models might need general scaffolding around properties like robustness, interpretability, efficiency, and privacy, before they are useful as vertical-specific systems—such problems are considered vertical-agnostic properties. While improvements along these aspects generally benefit many verticals, nonetheless, some of these aspects may also require vertical-specific considerations. Vertical adaptation: Designed for delivering specific value within verticals such as healthcare, web safety, and education, vertical-specific adaptations involve integrating large models’ capabilities with vertical-specific data, modeling, evaluation, and interfacing the capabilities with end-users of the vertical. Vertical-user intermediaries: These address vertical-agnostic challenges in interfacing systems built with large models with the end-users, focusing on aspects like trust calibration, feedback loops, and dynamic interfaces. While some interfacing challenges could be vertical-specific, others are broadly applicable challenges.
@@
3 Applying the framework: case studies
@@
Large AI models: The architectural scale and the pre-training dataset size have resulted in remarkable off-the-shelf capabilities of the large AI models, measured by their success on continually evolving benchmarks[12, 23]. While there has been a push towards even larger models trained on huge datasets, equipping them with the ability to support inputs in multiple modalities — language, vision, audio, and in some cases even sensor data [24], has the potential to unlock new applications across many verticals. For instance, multimodal models could, in principle, process radiology scans along with diagnostic questions [25], raw electrocardiogram (ECG) signals for health data analytics [26], provide voice-based tutoring [27], and write and review lengthy codebases [28]. While the development of these AI models offers an affordance for multimodal input, questions remain around how well the multimodal LLMs can reason over the non-textual forms of data[29, 30], which requires further work within this layer of the framework. Advances that ensure multimodal LLMs indeed model all modalities reliably will ensure greater off-the-shelf capabilities in the future iterations of these models.
@@
Vertical-agnostic properties: As we consider applying large AI models, the first set of problems are vertical-agnostic properties that apply to most verticals and pave the path for effective consideration of vertical-specific aspects. For instance, are multimodal LLMs robust to the plethora of plausible and realistic variations in user-provided inputs, given that it is unreasonable to assume users will constrain their inputs to the margins of the training distribution [16, 15, 31]? Will these models handle the personally identifiable information (PII) already encoded from its pre-training corpus [32] and the sensitive data provided by users while using it for iterative refinement (e.g., from patients [33] or students [34])? Relatedly, can large AI models provide interpretable predictions that foster transparency [35]? Addressing these problems as foundational steps will overcome bottlenecks across many verticals and enable a more effective use of the remedial techniques. Additionally, addressing these problems in the proximity of the large AI models layer could lead to integration of the remedial techniques in the development of upcoming large AI models; we elaborate on this in Section 4. We now move on to situating vertical-specific innovations, opportunities, and challenges within the framework, starting with healthcare and then exploring education. We chose to focus on healthcare and education verticals because they represent high-stakes domains with distinct yet complementary challenges: both verticals focus on building user-centered systems with healthcare demanding precision and rigorous safety measures while education prioritizing engagement and personalization to cater to diverse learner needs.
Vertical-User IntermediariesTrust calibration : How to best communicate models’ capabilities and limitations?Feedback loops : Capturing user feedback for iterative refinement of modelsDynamic interfaces : Design interfaces that engage users with higher and lower NFC levels
@@
Data: Curating data for tuning with medical dialogues Data: Curating data that supports better pedagogy
Modeling: Enabling patient history-taking for diagnoses Modeling: Enabling efficient learner-tuned pedagogy
Evaluation: Evaluating data quality by clinical standards Evaluation: Evaluating engagement and motivation
Interfacing: Clinician-AI collaboration for error correction Interfacing: Grounding responses in what the learner “sees”
Vertical-Agnostic PropertiesRobustness : Are models robust to realistic variations in input?Privacy : Do models carefully handle sensitive data/PII?Interpretability : Can models provide interpretable predictions?
@@
How can modalities beyond language (visual, audio, sensor data) be reliably processed?
@@
Table 1: Operationalizing the framework for adoption of large AI models in the healthcare and education verticals; read from bottom to top. Each question is an example of existing/potential innovations, opportunities, and challenges in vertical adoption of large artificial intelligence models.
@@
Vertical adaptation: (a) MLLMs for healthcare: One of the well-explored clinical applications of multimodal deep learning models, including the recent multimodal LLMs, is as an assistant for radiologists. The models can help clinicians make diagnoses via conversational-assistance as well as writing medical reports [36, 37]. Even though off-the-shelf LLMs encode medical knowledge [38], they need to be vertically adapted to acquire diagnostically useful information from natural conversations with primary care providers. This vertical adaptation involves curating the right data of diagnostic conversations, which would require close involvement of domain-experts [39]. Beyond curating the right data, off-the-shelf LLMs do not demonstrate properties that are required to accurately model the data–for instance, history-taking [40, 39]. In the context of radiology, this would involve equipping the underlying multimodal LLMs with temporal modeling capabilities [41]. To evaluate whether the modeling approach is effective on the curated data, vertical-agnostic evaluations may not suffice. For instance, Yu et al. (2023) [42] demonstrate that generic ‘natural language generation’ metrics are not effective in capturing clinically pertinent differences between AI-generated radiology reports and those written by experts, and propose meaningful metrics to guide future research in this vertical. The evaluations uncovered that while both experts and AI systems can make mistakes while generating radiology reports alone, the instances of inaccuracies decrease when experts and AI work in collaboration to fix the errors [40]. Nonetheless, a crucial challenge remains unaddressed when interfacing the AI systems with clinicians – the collaboration loses effectiveness when the expert either overly relies on the AI predictions [43, 44] or is excessively critical of them [45].
@@
Vertical adaptation: (b) MLLMs for Education: Large AI models, including multimodal LLMs, are already being used by students across the globe to assist with their learning [46], demonstrating their promise in tutoring. Even though their ad hoc capabilities are noteworthy, Macina et al. (2023)  [47] note that their systematic impact on “tutoring has largely remained unaffected”. A central challenge in this vertical concerns curating data that captures diverse pedagogical strategies, covering a broad range of topics, learner demographics, and instructional modes encountered in real classrooms [48]. From a modeling perspective, besides fine-tuning (which may prove to be an inefficient strategy for adapting to different definitions of what constitutes effective pedagogy), it is advantageous to allow learners to specify desired attributes across pedagogical dimensions and have the model reflect them [49]. The evaluation of such models should be grounded in learning science, which often prioritizes motivating and promoting engagement from the learner, and not just “giving the right answer”[50]. It is also crucial that the interfaces support learner-tutor interactions such that the conversations are grounded in what the student sees, which is again one of the core principles of student-centric pedagogy.
@@
Vertical-user intermediaries: Even though there are vertical-specific considerations involved in interfacing systems with users, some of these challenges are frequently encountered across many verticals. For instance, as we noted, clinician-AI collaboration tends to become less useful when the capabilities and limitations are not adequately understood by the experts. Similar observations around ‘algorithm aversion or appreciation’ exist across many other verticals [51, 52, 53]. Given the broad nature of the underlying challenge, it is effective to work on trust calibration as a vertical-agnostic vertical-user intermediary. For instance, introducing uncertainty expressions (“I’m not sure, but…”) in AI-generated responses leads to decreased over-reliance and calibrated trust among users [54] – an insight that can benefit many verticals, including generative information retrieval, healthcare assistance, and educational tutoring. Similarly, designing interfaces that capture real-time, in-situ, and implicit user feedback, will unlock iterative refinement of underlying systems across many verticals[55]. Furthermore, interfaces that dynamically adapt to deliver the optimal experience to a group of users with diverse Need For Cognition levels (NFC; a personality trait that considers users’ tendency to engage with cognitive activities[56]) is an opportunity that will benefit many verticals. More specifically, for decades, researchers have observed that users with higher NFC levels benefit significantly from complex interfaces, while others struggle to adopt them[57] — a pattern that spans many verticals and has continued with recent Generative AI technologies [58, 19].
4 Dynamism between layers of the framework
@@
Let us start with vertical-specific considerations influencing vertical-agnostic properties. As practitioners and researchers explore the applications of large models in different verticals, they highlight the successes as well as shortcomings. Certain aspects such as lack of robustness, poor handling of private data, and lack of efficient adaptability are frequently encountered/identified across verticals and make their way to being more effectively addressed in a vertical-agnostic manner. When addressing these challenges in a vertical-agnostic manner, vertical-agnostic solutions can influence the development of the next iterations of large AI models. Matryoshka representations [59] present a strong case study for this. In many real-world search systems (e.g., patient records, legal reviews, and educational video search), relevant items are often retrieved by computing similarity between neural representations. At large scales—hundreds of millions of items—this must be efficient. Kusupati et al. (2022) [59] introduced “Matryoshka doll” representations, where $14\times$14 × smaller embeddings match the performance of full-size ones for classification and retrieval. This innovation supports flexible representation sizes, enabling efficient retrieval and classification in a vertical-agnostic manner, even though the initial inadequacies were identified during many vertical-specific adaptations. Although originally the effectiveness of Matryoshka representations was demonstrated on ImageNet-scale datasets, it was later adopted by OpenAI and others to train large AI models (e.g., text-embedding-3) on web-scale data [60].
@@
Akin to the dynamism between vertical-specific considerations, vertical-agnostic properties, and large AI models, certain challenges may also shift from interfacing-related aspects within verticals to broader vertical-user intermediaries. Consider AI hallucinations in high-stakes fields like healthcare, where unreliable responses can lead to adverse outcomes and erode trust, prompting research on cognitive forcing interventions in medical AI-assisted decision-making [19]. With LLMs now prevalent in many domains (e.g., web retrieval, law, education, and mental healthcare), hallucination has become a key challenge. This has spurred work on vertical-user intermediaries, including quantifiable or language-based uncertainty cues [61, 54] and referencing source documents for verifiable claims [62], with early evidence of reduced user over-reliance on AI outputs [63].
@@
The dynamism between layers is inherently healthy as it facilitates the distributed-yet-collaborative development of AI systems. The large AI models are often developed in highly resource-rich environments, while the practitioners and researchers who develop vertical-specific insights that trickle down to the bottom layers of the framework are often situated outside of these selectively few environments. In this context, two points are worth noting: first, to deliver benefits within a vertical using large AI models, it is critical to actively engage with domain experts who have curated specialized data, engineered novel methods, and designed meaningful evaluations for those verticals. This enables specialized insights that arose from vertical-specific explorations, while simultaneously embedding them in more general-purpose model architectures. Second, it is incorrect to assume that vertical-specific explorations merely piggyback on advances in large AI models. Rather, vertical-specific advances play a critical role in the development of large AI models. When used carefully, these vertical-specific insights can make general-purpose models more capable and speed up the release of solutions for specific needs among users.
5 Intended framework outcomes
@@
5.1 Optimally situating and borrowing innovations

### v2

@@ Introduction
Large artificial intelligence (AI) models have long served as powerful tools for advancing domain-specific research and development. Early examples include the adaptation of language embeddings (like word2vec (Mikolov et al. 2013)) for generating disease taxonomies (Ghosh et al. 2016), as well as the use of YOLO-based image models  (Redmon 2016) for conducting animal population censuses  (Parham et al. 2017). More recently, large models such as GPT-4o  (Hurst et al. 2024) and SAM  (Kirillov et al. 2023) have demonstrated more advanced capabilities, catalyzing further innovations in domains as varied as AI tutoring  (Lin, Huang, and Lu 2023) and digital pathology  (Deng et al. 2023), all the way to software development  (Barenkamp, Rebstadt, and Thomas 2020). The success of these models has been propelled by the increasing scale of their architectures—rising from a few million parameters in 2013, to hundreds of millions in 2018, and now surpassing a trillion parameters in the most recent deployments (Elmeleegy et al. 2024)—as well as by the sheer volume of data used for their training (Bahri et al. 2024). Many widely used benchmarks, both general-purpose and domain-specific, have reached saturation or are rapidly approaching it  (Chollet 2024; Phan et al. 2025). At the same time, frameworks built around these large models (such as Application Programming Interfaces; APIs) have lowered barriers to entry (Schillaci 2024), spurring a surge in applications and attracting researchers and practitioners from a broad range of disciplines.
@@
Despite these successes, a decade of adapting large models in diverse verticals has highlighted persistent challenges. For example, research teams have discovered that BERT and CLIP-based models can be brittle to small input variations  (Verma et al. 2022b; Ramshetty, Verma, and Kumar 2023), large language models (LLMs) often exhibit sensitivity to prompt formatting  (Sclar et al. 2024), and performance of large models can diminish in highly specialized settings (Deng et al. 2023; Chandra et al. 2024). Moreover, AI systems sometimes struggle to effectively support diverse user groups, such as users with a lower Need For Cognition  (Buçinca, Malaya, and Gajos 2021). These challenges, taken together, pose bottlenecks in the vertical-adoption of the large models. The process of adapting large models to verticals demands coordinated efforts from stakeholders across varied disciplines and continues to be an active area of research. Existing frameworks provide good starting points to address some of these problems at a more granular level—either focusing on one component of developing AI systems or a specific vertical. For instance, Ehsan et al. (2023) present a framework to bridge the gap between social and technical aspects of developing explainable AI systems. On the other hand, focusing on a specific vertical, Trotsyuk et al. (2024) present a framework to address potential misuse of AI in biomedical research. There is still an unmet need for a more comprehensive framework comprising various components required for vertical adoption of large models and has broad applicability towards generalizing to many verticals.
@@
Our proposed framework consists of 4 layers (see Figure 1), (i) starting with large AI models at the bottom, (ii) vertical-agnostic properties, (iii) vertical adaptation, and (iv) finally, vertical-user intermediaries. The 4 layers represent step-wise modular abstractions involved in developing systems with large models that deliver practical value to their intended users. Drawing on case studies from multiple verticals, we discuss how to situate innovations, challenges, and opportunities within each layer of this framework. We consider innovations as advances in algorithms, metrics, or interface designs that resolve identified pain‑points; challenges as specific obstacles that hinder vertical adoption of large AI models; and opportunities as recurring unmet needs that signal room for high‑leverage solutions. We use the term situating to indicate anchoring the contributions in one of the four layers of the framework. The overall aim of the situating contributions within the framework is to ensure that the holistic adoption of large AI models across various verticals remains effective.
Beyond describing the framework and its layers, we also highlight the inherent dynamism among these layers (i.e., how they interact with and influence one another over time). Finally, drawing on observed trends, we present actionable recommendations that benefit researchers and developers across various verticals and disciplines. For instance, we discuss whether aspects like robustness and privacy are better addressed as vertical-agnostic properties or as a vertical-specific concerns, how scoping feedback from many verticals and intermediary problems can lead to more effective development of newer large models, when domain-specific experts can borrow modeling and interfacing techniques from others while innovating on the domain-specific data curation and evaluation methods, and how interfacing AI systems with users remains ripe with opportunities. Collectively, we believe that adopting our framework will (a) guide optimally placed innovations, (b) highlight potential opportunities, and (c) enable cross-disciplinary dialogue.
Who can benefit from the framework? The framework is useful for interdisciplinary teams who want to adopt large AI models in their respective verticals, and for researchers who hope to position their work within a broad ecosystem for identification of cross-layer connections and translation of knowledge across contexts. For teams focused on a single vertical, the framework decomposes the adoption pipeline into modular components. Across multiple verticals, it provides a holistic view of the broader ecosystem that (a) fosters cross-vertical exchange of innovations, opportunities, and challenges (i.e., what can teams in healthcare learn from teams in education) as well as (b) funnels feedback from many verticals to improve the next iteration of large artificial intelligence models.
A Framework for Advancing Vertical Systems With AI Models
@@
1. Large AI models: These are large-scale (in terms of training dataset size and number of parameters) AI models, more recently dubbed foundation models (Bommasani et al. 2021) that power diverse applications across verticals. These include modality-specific models (e.g., language-only, vision-only) and multimodal models capable of handling multiple input modalities. Their functional utility lies in the general off-the-shelf capabilities they provide (e.g., natural language understanding, image-text alignment) and adaptability to new verticals through techniques such as fine-tuning, prompting, or in-context learning.
2. Vertical-agnostic properties: Large models might need general scaffolding around properties like robustness, interpretability, efficiency, and privacy, before they are useful as vertical-specific systems—such problems are considered vertical-agnostic properties. While improvements along these aspects generally benefit many verticals, nonetheless, some of these aspects may also require vertical-specific considerations.
3. Vertical adaptation: Designed for delivering specific value within verticals such as healthcare, web safety, and education, vertical-specific adaptations involve integrating large models’ capabilities with vertical-specific data, modeling, evaluation, and interfacing the capabilities with end-users of the vertical.
4. Vertical-user intermediaries: These address vertical-agnostic challenges in interfacing systems built with large models with the end-users, focusing on aspects like trust calibration, feedback loops, and dynamic interfaces. While some interfacing challenges could be vertical-specific, others are broadly applicable challenges.
@@
Applying the Framework: Case Studies
@@
Large AI models: The architectural scale and the pre-training dataset size have resulted in remarkable off-the-shelf capabilities of the large AI models, measured by their success on continually evolving benchmarks(Chollet 2024; Liang et al. 022). While there has been a push towards even larger models trained on huge datasets, equipping them with the ability to support inputs in multiple modalities — language, vision, audio, and in some cases even sensor data (Moon et al. 2024), has the potential to unlock new applications across many verticals. For instance, multimodal models could, in principle, process radiology scans along with diagnostic questions (Bhayana 2024), raw electrocardiogram (ECG) signals for health data analytics (Quer and Topol 2024), provide voice-based tutoring (Katsarou et al. 2023), and write and review lengthy codebases (Bairi et al. 2024). While the development of these AI models offers an affordance for multimodal input, questions remain around how well the multimodal LLMs can reason over the non-textual forms of data(Tong et al. 2024; Verma et al. 2024a), which requires further work within this layer of the framework. Advances that ensure multimodal LLMs indeed model all modalities reliably will ensure greater off-the-shelf capabilities in the future iterations of these models.
@@
Vertical-agnostic properties: As we consider applying large AI models, the first set of problems are vertical-agnostic properties that apply to most verticals and pave the path for effective consideration of vertical-specific aspects. For instance, are multimodal LLMs robust to the plethora of plausible and realistic variations in user-provided inputs, given that it is unreasonable to assume users will constrain their inputs to the margins of the training distribution (Ramshetty, Verma, and Kumar 2023; Verma et al. 2022b; Nookala et al. 2023)? Will these models handle the personally identifiable information (PII) already encoded from its pre-training corpus (Carlini et al. 2021) and the sensitive data provided by users while using it for iterative refinement (e.g., from patients (Pan et al. 2024) or students (Yang and Beil 2024))? Relatedly, can large AI models provide interpretable predictions that foster transparency (Stiglic et al. 2020)? Addressing these problems as foundational steps will overcome bottlenecks across many verticals and enable a more effective use of the remedial techniques. Additionally, addressing these problems in the proximity of the large AI models layer could lead to integration of the remedial techniques in the development of upcoming large AI models; we elaborate on this in Section Dynamism Between Layers of the Framework. We now move on to situating vertical-specific innovations, opportunities, and challenges within the framework, starting with healthcare and then exploring education. We chose to focus on healthcare and education verticals because they represent high-stakes domains with distinct yet complementary challenges: both verticals focus on building user-centered systems with healthcare demanding precision and rigorous safety measures while education prioritizing engagement and personalization to cater to diverse pedagogical needs of the learner.
Vertical-User IntermediariesTrust calibration : How can AI systems effectively communicate their capabilities and limitations to users to prevent issues of overreliance or unwarranted skepticism?Feedback loops :How can real-time user feedback be systematically captured and integrated into iterative refinement processes for large AI models?Dynamic interfaces : How can interface designs dynamically accommodate users with varying cognitive engagement preferences, optimizing usability for both high and low Need For Cognition (NFC) individuals?
@@
Data: What methodologies can be used to curate specialized datasets of medical dialogues necessary for accurately tuning multimodal models for clinical use? Data: How should educational data be curated to reflect diverse pedagogical strategies and learner demographics to improve AI-assisted tutoring?
Modeling: How can AI models effectively incorporate temporal patient history data to improve accuracy in clinical diagnostics? Modeling: What modeling techniques can efficiently adapt AI systems to individual learner needs and preferred pedagogical strategies?
Evaluation: Which clinical standards and metrics most effectively measure the quality and reliability of AI-generated outputs in healthcare settings? Evaluation: How can AI system evaluations accurately assess their impact on learner motivation, engagement, and educational outcomes beyond correctness of content?
Interfacing: How can AI systems best support collaboration with clinicians? Interfacing: How can AI tutoring systems effectively ground responses within the immediate learning context and visual perspective of the student?
Vertical-Agnostic PropertiesRobustness : How robust are AI models to realistic & plausible variations in user inputs spanning multiple modalities?Privacy : How effectively can AI models ensure the privacy & security of personal and sensitive data?Interpretability : In what ways can AI models be made to provide interpretable predictions to improve transparency?
@@
How can modalities beyond language (visual, audio, sensor data) be reliably processed with large AI models?
@@
Table 1: Operationalizing the framework for adoption of large AI models in the healthcare and education verticals; read from bottom to top. Each question is an example of existing/potential innovations, challenges, and opportunities in vertical adoption of large artificial intelligence models (in this case, multimodal large language models).
@@
Vertical adaptation: (a) MLLMs for healthcare: One of the well-explored clinical applications of multimodal deep learning models, including the recent multimodal LLMs, is as an assistant for radiologists. The models can help clinicians make diagnoses via conversational-assistance as well as writing medical reports (Johri et al. 2025; Zhang et al. 2024). Even though off-the-shelf LLMs encode medical knowledge (Singhal et al. 2023), they need to be vertically adapted to acquire diagnostically useful information from natural conversations with primary care providers. This vertical adaptation involves curating the right data of diagnostic conversations, which would require close involvement of domain-experts (Tu et al. 2024). Beyond curating the right data, off-the-shelf LLMs do not demonstrate properties that are required to accurately model the data–for instance, history-taking (Tanno et al. 2024; Tu et al. 2024). In the context of radiology, this would involve equipping the underlying multimodal LLMs with temporal modeling capabilities (Bannur et al. 2023). To evaluate whether the modeling approach is effective on the curated data, vertical-agnostic evaluations may not suffice. For instance, Yu et al. (2023) (Yu et al. 2023) demonstrate that generic ‘natural language generation’ metrics are not effective in capturing clinically pertinent differences between AI-generated radiology reports and those written by experts, and propose meaningful metrics to guide future research in this vertical. The evaluations uncovered that while both experts and AI systems can make mistakes while generating radiology reports alone, the instances of inaccuracies decrease when experts and AI work in collaboration to fix the errors (Tanno et al. 2024). Nonetheless, a crucial challenge remains unaddressed when interfacing the AI systems with clinicians – the collaboration loses effectiveness when the expert either overly relies on the AI predictions (Seah et al. 2021; Rajpurkar et al. 2020) or is excessively critical of them (Agarwal et al. 2023).
@@
Vertical adaptation: (b) MLLMs for Education: Large AI models, including multimodal LLMs, are already being used by students across the globe to assist with their learning (Zhu, Zhang, and Wang 2024), demonstrating their promise in tutoring. Even though their ad hoc capabilities are noteworthy, Macina et al. (2023)  (Macina et al. 2023) note that their systematic impact on “tutoring has largely remained unaffected”. A central challenge in this vertical concerns curating data that captures diverse pedagogical strategies, covering a broad range of topics, learner demographics, and instructional modes encountered in real classrooms (Jurenka et al. 2024). From a modeling perspective, besides fine-tuning (which may prove to be an inefficient strategy for adapting to different definitions of what constitutes effective pedagogy), it is advantageous to allow learners to specify desired attributes across pedagogical dimensions and have the model reflect them (Team et al. 2024). The evaluation of such models should be grounded in learning science, which often prioritizes motivating and promoting engagement from the learner, and not just “giving the right answer”(Foster et al. 2023). It is also crucial that the interfaces support learner-tutor interactions such that the conversations are grounded in what the student “sees”, which is one of the core principles of student-centric pedagogy.
@@
Vertical-user intermediaries: Even though there are vertical-specific considerations involved in interfacing systems with users, some of these challenges are frequently encountered across many verticals. For instance, as we noted, clinician-AI collaboration tends to become less useful when the capabilities and limitations are not adequately understood by the experts. Similar observations around ‘algorithm aversion or appreciation’ exist across other verticals (Dietvorst, Simmons, and Massey 2015; Logg, Minson, and Moore 2019; Qian and Wexler 2024). Given the broad nature of the underlying challenge, it is effective to work on trust calibration as a vertical-agnostic vertical-user intermediary. E.g., introducing uncertainty expressions (“I’m not sure, but…”) in AI-generated responses leads to decreased over-reliance and calibrated trust among users (Kim et al. 2024) – an insight that can benefit many verticals, including generative information retrieval, healthcare assistance, and educational tutoring. Similarly, designing interfaces that capture real-time, in-situ, and implicit user feedback, will unlock iterative refinement of underlying systems across many verticals(Shi et al. 2024). Furthermore, interfaces that dynamically adapt to deliver the optimal experience to a group of users with diverse Need For Cognition levels (NFC; a personality trait that considers users’ tendency to engage with cognitive activities(Cacioppo and Petty 1982)) is an opportunity that will benefit many verticals. More specifically, for decades, researchers have observed that users with higher NFC levels benefit significantly from complex interfaces, while others struggle to adopt them(Carenini 2001) — a pattern that spans many verticals and has continued with recent generative artificial intelligence technologies (Toner-Rodgers 2024; Buçinca, Malaya, and Gajos 2021).
Dynamism Between Layers of the Framework
@@
Let us start with vertical-specific considerations influencing vertical-agnostic properties. As practitioners and researchers explore the applications of large models in different verticals, they highlight the successes as well as shortcomings. Certain aspects such as lack of robustness, poor handling of private data, and lack of efficient adaptability are frequently encountered/identified across verticals and make their way to being more effectively addressed in a vertical-agnostic manner. When addressing these challenges in a vertical-agnostic manner, vertical-agnostic solutions can influence the development of the next iterations of large AI models. Matryoshka representations (Kusupati et al. 2022) present a strong case study for this. In many real-world search systems (e.g., patient records, legal reviews, and educational video search), relevant items are often retrieved by computing similarity between neural representations. At large scales—hundreds of millions of items—this must be efficient. Kusupati et al. (2022) (Kusupati et al. 2022) introduced “Matryoshka doll” representations, where $14\times$ smaller embeddings match the performance of full-size ones for classification and retrieval. This innovation supports flexible representation sizes, enabling efficient retrieval and classification in a vertical-agnostic manner, even though the initial inadequacies were identified during many vertical-specific adaptations. Although originally the effectiveness of Matryoshka representations was demonstrated on ImageNet-scale datasets, it was later adopted by OpenAI and others to train large AI models (e.g., text-embedding-3) on web-scale data (OpenAI 2024).
@@
Akin to the dynamism between vertical-specific considerations, vertical-agnostic properties, and large AI models, certain challenges may also shift from interfacing-related aspects within verticals to broader vertical-user intermediaries. Consider AI hallucinations in high-stakes fields like healthcare, where unreliable responses can lead to adverse outcomes and erode trust, prompting research on cognitive forcing interventions in medical AI-assisted decision-making (Buçinca, Malaya, and Gajos 2021). With LLMs now prevalent in many domains (e.g., web retrieval, law, education, and mental healthcare), hallucination has become a key challenge. This has spurred work on vertical-user intermediaries, including quantifiable or language-based uncertainty cues (Xiong et al. 2024; Kim et al. 2024) and referencing source documents for verifiable claims (Gao et al. 2023), with early evidence of reduced user over-reliance on AI outputs (Bo, Wan, and Anderson 2024).
@@
The dynamism between layers is inherently desirable as it facilitates the distributed-yet-collaborative development of AI systems. The large AI models are often developed in highly resource-rich environments, while the practitioners and researchers who develop vertical-specific insights that trickle down to the bottom layers of the framework are often situated outside of these selective environments. Here, two points are worth noting: first, to deliver benefits within a vertical using large AI models, it is critical to actively engage with domain experts who have curated specialized data, engineered novel methods, or designed meaningful evaluations for those verticals. This enables specialized insights that arose from vertical-specific explorations, while simultaneously embedding them in more general-purpose model architectures. Second, it is incorrect to assume that vertical-specific explorations merely piggyback on advances in large AI models. Rather, vertical-specific advances play a critical role in the development of large AI models. When used carefully, these vertical-specific insights can make general-purpose models more capable and speed up the release of solutions for specific user needs.
Intended Framework Outcomes
@@
Optimally situating and borrowing innovations